OpenStudy (anonymous):

r=1/3d(60-t) solve for t.

OpenStudy (danjs):

hello, welcome to the site

OpenStudy (danjs):

\[r = \frac{ 1 }{ 3d(60-t) }\] is that what you have?

OpenStudy (danjs):

\[r = \frac{ 1 }{ 3 }*d(60-t)\] or that

OpenStudy (anonymous):

yes the second one

OpenStudy (anonymous):

\[r=\frac{ 1 }{ 3}*d(60-t)\]

OpenStudy (anonymous):

solve for T. whats the first step?

OpenStudy (danjs):

k

OpenStudy (danjs):

first multiply both sides of it by 3 to clear out the fraction

OpenStudy (danjs):

what does that give ya

OpenStudy (anonymous):

\[3*r=1*d(60-t)\] ?

OpenStudy (danjs):

hi, are you still here?

OpenStudy (danjs):

That is the right first step

OpenStudy (danjs):

in order to solve for t, we need to distribute the d into the parenthesis term first 3r = d*(60 - t) 3r = 60*d - d*t

OpenStudy (danjs):

The distribution looks like this: